OpenStudy (anonymous):

C2H2 + 2H2 mc022-1.jpg C2H6 How many moles of hydrogen are needed to produce 13.78 mol of ethane?

OpenStudy (anonymous):

\[\frac{ mol of ethane }{ 1 }\times \frac{ moles of H2}{ mol of ethane }\]

OpenStudy (anonymous):

Without numbers can you figure it out?